OpenStudy (90890908967777878888):

ken swam for 2 1/2 on friday and 1 11/12 hours on Saturday what is the total amount he swim for these days

OpenStudy (90890908967777878888):

@ChantySquirrel1129**

OpenStudy (anonymous):

\[2+\frac{1}{2}+1+\frac{11}{12}=\frac{53}{12}=4+\frac{5}{12} \]I assume that you know how to deal with fractions.

OpenStudy (anonymous):

5/12 hours is 25 minutes.

OpenStudy (90890908967777878888):

yea